### v1

@@ Abstract
Recently, a new general wave phenomenon, namely “the anti-localization of non-stationary linear waves”, has been introduced and discussed. This is zeroing of the propagating component for a non-stationary wave-field near a defect in infinitely long wave-guides. The phenomenon is known to be observed in both continuum and discrete mechanical systems with a defect, provided that the frequency spectrum for the corresponding homogeneous system possesses a stop-band. In this paper, we show that the anti-localization is also quite common for nonlinear systems. To demonstrate this, we numerically solve several non-stationary problems for an infinite $\beta$-FPUT chain with a defect.
1 Introduction
@@
In infinite uniform linear continuum and discrete systems, the perturbations caused by a pulse point load travel as waves [^1] For a discrete system, it is more correct to speak about quasi-waves. along a system from the load application point to infinity. The presence of an inhomogeneity (or a defect) at the load application point can lead to accumulation or decumulation of the wave energy near the defect compared to the uniform case. Apparently, the first study where this fact was demonstrated is the classical paper by Lamb Lamb1900 , where it was shown that embedding a discrete mass-spring system into a continuum medium described by the 1D wave equation leads to the leakage of energy of a mass-spring system. On the other hand, such an embedding leads to the accumulation of energy in the wave-guide near the inhomogeneity, since in a uniform medium, described by the 1D wave equation, the perturbations caused by a point pulse load immediately leave the neighbourhood of the application point, leaving it in the rest [^2] This is due to the structure of the fundamental solution for the 1D wave equation Vladimirov1971 ; Watanabe2015 ..
@@
In uniform linear dispersive systems, the amplitude of perturbations caused by a point pulse load vanishes with time in a continuous manner. Generally, in 1D systems the order of decay for perturbations (e.g., for the particle velocities, the strains, or the displacements if the system is fixed somewhere and does not have rigid modes of motion) is $1/\sqrt{t}$, where $t\to\infty$ is time. This result is well-known; see, e.g., Whitham1999 ; Slepyan1972 , and can be obtained by the asymptotic method of stationary phase applied to a representation of the exact non-stationary solution in the form of a Fourier integral. Again, introducing the inhomogeneity can lead to the accumulation or decumulation of the wave energy nearby.
@@
The effect of the accumulation can be related to the so-called trapped modes ursell1951trapping ; Ind-book-R2E ; Kuznetsov2002 ; Andrianov2013 ; Gavrilov2024nody , which can exist in dispersive systems. In infinite discrete systems (lattices or chains), the analogous phenomenon is known as the impurity-induced localized modes Montroll1955 ; Kivshar1991 ; Kivshar1997 ; Mamalui1999 . The existence of such modes leads to trapping of some portions of the wave energy of a non-stationary process forever near a load application point at an inhomogeneity kaplunov1986torsional ; Teramoto1960 (we call this the localization of non-stationary waves). The whole frequency spectrum of our system can have a stop-band, where complex wave-numbers correspond to real frequencies. In such systems, so-called strong type of localization luongo2001mode ; Luongo1992 is quite common, which is related to a pole in the denominator of the system Green function in the frequency domain lying inside the stop-band. The rest of the energy (or all the wave energy if there is no localization) still travels to infinity as a vanishing propagating component of the wave-field, where perturbations again have the order of decay $1/\sqrt{t}$ (in a 1D system).
@@
Recently Shishkina2023cmat ; Shishkina2023jsv ; Gavrilov2024nody , an opposite general wave phenomenon, which has been called the anti-localization of non-stationary waves, related to the decumulation of the energy near an inhomogeneity, has been introduced. It is observed in both discrete and continuum infinite linear systems, having the dispersion graph with a stop-band [^3] The frequency spectrum of a discrete linear infinite system always has a stop-band Kossevich1999 for high values of the frequency.. The presence of a stop-band leads to the existing of the cut-off frequency to which zero group velocity corresponds. The oscillation corresponding to the cut-off frequency accumulates near a point pulse source hemmer1959dynamic ; slepyan1987energy ; this can cause the phenomenon of the anti-localization. The existence of such a phenomenon is in agreement with some particular results previously obtained kaplunov1986torsional ; Mueller1962 ; Mueller2012 ; Kashiwamura1962 ; Rubin1963 ; hemmer1959dynamic for some systems, where the strong localization can take place. The anti-localization is zeroing of the leading-order term of order $1/\sqrt{t}$ in asymptotic expansions for perturbations at an inhomogeneity. Consequently, the next term of order $1/\sqrt{t^{3}}$ becomes a leading-order term there. Accordingly, the perturbations, which correspond to the propagating component, in an expanding with time neighbourhood of the inhomogeneity, become asymptotically small compared to ones in zones closer to the leading wave-fronts. The amplitude of oscillation in the propagating component has a clear almost zero minimum at the inhomogeneity. Note that the anti-localization can be observed in the absence of the localization as well as co-exist with the localization. In the latter case, after some time, a peak of amplitude for oscillation also becomes clearly recognizable, being observed in the centre of a zone where perturbation amplitude is almost zero.
@@
In the paper, we show that the anti-localization is not a specific peculiarity of linear systems. We demonstrate that it is also common for nonlinear ones. To achieve this, we solve several non-stationary problems for an infinitely long $\beta$-FPUT chain, introduced in the famous study Fermi1955 , with a defect. In the nonlinear case, we have no possibility to investigate the anti-localization in an analytic way using the method of stationary phase. Accordingly, we use a pure numerical approach, considering the previously obtained solutions for the corresponding linear systems as the reference solutions. All numerical calculations presented in the paper have been performed twice using completely independent code and different numerical approaches. The first approach is to use routine scipy.integrate.odeint included in python framework scipy scipy , which is actually based on lsoda from the fortran library odepack. The second approach is related to the self-written code utilizing the semi-implicit Euler method Skeel2020 . The results obtained by both approaches seem almost identical. The choice of a discrete system (not a continuum one) as the system under investigation is caused by the considerations of the simplicity of numerical solving.
@@
In our opinion, the anti-localization of non-stationary waves essentially influences the processes of energy (heat) transfer in linear and nonlinear lattices. In particular, the emergence of cold points at the defects demonstrated in linear and nonlinear cases Gendelman2021 , apparently, can be associated with this phenomenon, though the mathematical formulation used in Gendelman2021 is different. The existence of the anti-localization makes the single-mode approximation very effective Gavrilov2024nody when looking for approximate non-stationary solutions describing non-stationary oscillation in the systems where the strong localization occurs.
@@
The structure of the paper is as follows. In Sect. 2, we discuss the mathematical formulation of the problem in the dimensionless form (the non-dimensionalization is provided in Appendix A). In Sects. 36, several nonlinear problems, where we expect the emergence of the anti-localization, are investigated. In Conclusion (Sect. 7), we discuss the basic results of the paper.
2 The mathematical formulation
@@
Consider an infinite ordered chain of point particles with nearest-neighbour interactions. All particles have an equal mass, whereas one particle has an alternated mass. The particles are labeled by the discrete variable $n\in\mathbb{Z}$, where $\mathbb{Z}$ is the set of all integers. Without loss of generality, we assume that the alternated mass corresponds to $n=0$. The particles are connected by nonlinear springs (i.e., by the bonds) with cubic nonlinearity also labeled by $n\in\mathbb{Z}$. We assume that a spring label equals the label of the preceding particle. All springs have equal stiffness excepting two of them with numbers $n=-1$ and $n=0$, which connect the particle with $n=0$ and its neighbours. Thus, we consider a $\beta$-FPUT chain with a point impurity (or a defect) in the form of a mass-spring inclusion. Everywhere through the paper, excepting Sect. 6, we assume that the inclusion is symmetric [^4] In Sect. 6 we, in particular, consider a linear chain with an inclusion in the form of a nonlinear spring with the linear stiffness being equal to the spring stiffness in the rest of the chain.. The schematic of the system is presented in Fig. 1.
@@
Here, overdot denotes the derivative with respect to the dimensionless time $t$; $u_{n}(t)$ is the dimensionless displacement of the particle with a label $n\in\mathbb{Z}$; $\beta$ is the dimensionless cubic stiffness for the bonds outside the defect; $m>0$ is the dimensionless mass of the particle at the defect; $K$ and $B$ are the linear and cubic dimensionless stiffnesses for the bonds at the defect, respectively; $\delta_{n}$ is the Kronecker delta ($1$ if and only if $n=0$, 00 otherwise, $n\in\mathbb{Z}$). The initial conditions are:
$u_{n}(0)=0,\qquad v_{n}(0)=m_{N}^{-1}\delta_{n-N}.$ (2.6)
Initial conditions in the form of Eq. (2.6) correspond to the impulse force $\delta(t)$ applied to the particle with label $N$, see Appendix A. Here $\delta(t)$ is the Dirac delta-function.
Note that we consider the system with a finite nonlinearity, i.e., the dimensionless parameter
$\bar{\beta}=\max_{n}\big{(}|\beta_{n}|\big{)}$ (2.7)
is assumed to be of order one, i.e., not a small quantity. The finite nonlinearity makes linear and nonlinear results to be clearly different for values of time $t$ such that the leading-order term of the perturbation asymptotics calculated for the corresponding linear system by the method of stationary phase eventually becomes close to the numeric results for the linear system.
@@
$m=1,\qquad K=1,\qquad B_{L}=B_{R}=\beta=0$ (2.8)
of an infinite uniform linear chain, the exact solution of the problem defined by Eqs. (2.1)–(2.6) for the particle velocities $v_{n}(t)$ is $V_{n-N}$ schrodinger1914dynamik ; Hamilton1940 ; Havelock1910 ; Muehlich2020 :
$V_{n}=J_{2n}(2t)=J_{2|n|}(2t),$ (2.9)
@@
3 An infinite $\beta$-FPUT chain with an isotopic defect
@@
In the linear case $\beta=0$, this problem is considered in studies Rubin1963 ; Shishkina2023cmat . In Shishkina2023cmat an approximate solution of asymptotic nature, describing the wave-field as a whole and valid for large values of time $t$, is obtained. We reproduce here some important formulae from Shishkina2023cmat to explain to the reader what the anti-localization of non-stationary waves is. The exact solution for the particle velocities in the integral form is an inverse Fourier transform:
${v}_{n}=-\frac{\mathrm{i}}{2\pi}\left(\int_{\mathbb{P}}+\int_{\mathbb{S}}\right)\Omega\mathcal{G}_{n}(\Omega)\mathrm{e}^{-\mathrm{i}\Omega t}\,\mathrm{d}\Omega+\mathrm{c.c.}\\ =I_{n}^{\mathrm{pass}}+I_{n}^{\mathrm{stop}}+\mathrm{c.c.}=v_{n}^{\mathrm{pass}}+v_{n}^{\mathrm{stop}},$ (3.6)
$\mathcal{G}_{n}(\Omega)=-\frac{\mathrm{e}^{\mathrm{i}|n|\operatorname{sign}\Omega\,\arccos\frac{2-\Omega^{2}}{2}}}{(m-1)\Omega^{2}+\mathrm{i}\Omega\sqrt{4-\Omega^{2}}},\ \ \Omega\in\mathbb{P};$ (3.7)
$\mathcal{G}_{n}(\Omega)=\frac{(-1)^{|n|}2^{|n|}}{\Phi^{|n|-1}(\Omega)\big{(}(-m\Omega^{2}+2)\Phi(\Omega)+4\big{)}},\ \ \Omega\in\mathbb{S};$ (3.8)
$\mathrm{c.c.}$ are the complex conjugate terms. Here,
$\Phi(\Omega)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\Omega^{2}-2+|\Omega|\sqrt{\Omega^{2}-4};$ (3.9)
=$\mathbb{P}\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}[0,\Omega_{\ast}],\qquad\mathbb{S}\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}(\Omega_{\ast},\infty),$ (3.10)
Remark 3.1
In Shishkina2023cmat mostly the heat, i.e., the kinetic energy, transfer processes are under consideration. Thus, the solution is obtained for the particle velocities, not for the displacements. This is also useful because the displacements involve a non-vanishing non-oscillating component (a rigid body mode of motion).
$10<m<$ 1 . (3.11)
If the last inequality is true, in the interval $(0,\Omega_{\ast})$ there exists a simple root $\Omega_{0}$ of the denominator for $\mathcal{G}^{\mathrm{stop}}$ such that
$\Omega_{0}=\frac{2}{\sqrt{m(2-m)}},$ (3.12)
and, therefore, integral $I^{\mathrm{stop}}$ does not exist in the classical sense. In the latter case, integral $I^{\mathrm{stop}}$ should be considered as the Fourier transform for a generalized function. To regularize this, we can apply Shishkina2023jsv ; Shishkina2023cmat the limit absorption principle. Finally, for $t\to\infty$ at a fixed position $n$ one gets
$v_{n}^{\mathrm{stop}}=I^{\mathrm{stop}}+\mathrm{c.c.}=H(1-m)L_{n}(t)+O(t^{-1}),$ (3.13)
${L}_{n}=\frac{2(m-1)(-1)^{|n|+1}m^{|n|-1}H(1-m)}{(2-m)^{|n|+1}}\cos\Omega_{0}t,$ (3.14)
where $t\to\infty$, see Shishkina2023cmat . Here $L_{n}(t)$ is the localized non-vanishing oscillation, $H(\cdot)$ is the Heaviside step-function.
The phenomenon of the anti-localization is mostly related to the integral $I^{\mathrm{pass}}$, which describes the propagating part of the wave-field. Following to Gavrilov2022ijhmt , we have estimated it in Shishkina2023cmat on a moving at an arbitrary sub-critical speed $w$ point of observation. Taking into account that solution (3.6) clearly is an even function of $n$, put
$|n|=wt,$ (3.15)
$1\displaystyle 0<w<$ 1 (3.16)
in the expression for $I^{\mathrm{pass}}$ defined by Eq. (3.6). The obtained integral can be estimated for $t\to\infty$ using the method of stationary phase. This provides an asymptotic continuum solution:
$v_{n}^{\mathrm{pass}}=I^{\mathrm{pass}}+\mathrm{c.c.}\\ =\frac{{A}(w)}{\sqrt{t}}\cos\Big{(}\phi(w)\,t+\frac{\pi}{4}+\psi(w)\Big{)}+O(t^{-3/2}),$ (3.17)
$\begin{gathered}{A}(w)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\!\frac{H(1-w)\,w}{\sqrt{\pi}{(1-w^{2})^{1/4}}\big{(}(m-1)^{2}(1-w^{2})+w^{2}\big{)}^{1/2}},\\ \\ \end{gathered}$ (3.21)
$\phi(w)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}2(w\arccos w-\sqrt{1-w^{2}}),$ (3.22)
$\psi(w)\stackrel{{\scriptstyle\mbox{\scriptsize def}}}{{=}}\arctan\frac{(m-1)\sqrt{1-w^{2}}}{w}.$ (3.23)
Note that for $w>1$ there is no stationary point, and the corresponding term of order $1/\sqrt{t}$ is zero. This fact is taken into account by introducing the multiplier $H(1-w)$ in the numerator of the right-hand side of Eq. (3.21). In any system with an isotopic impurity, i.e., in the case (3.2), we have:
${A}(w)=\frac{\sqrt{2}w}{\sqrt{\pi}|m-1|}+O(w^{3}),\quad w\to 0;\qquad A(0)=0;$ (3.24)
and, thus, the amplitude $A(w)$ of the propagating part $v^{\mathrm{pass}}$ for the particle velocities is small in a certain expanding (since $w=|x|/t$) neighbourhood of the impurity. We call this phenomenon the anti-localization of non-stationary waves. Finally, the asymptotic expansion for the right-hand side of Eq. (3.6) at $n=0$ (just at the impurity) has the following form
$v_{0}(t)=H(M-K)L_{0}(t)+\frac{\cos\left(2t-\frac{3\pi}{4}\right)}{2\sqrt{\pi}(m-1)^{2}\,t^{3/2}}\\ +o(t^{-3/2}).$ (3.25)
The first term in the right-hand side of Eq. (3.25) is the contribution from the poles ±$\pm\Omega_{0}$ and describes localized oscillation, which exists if and only if $m<1$. The second term is the anti-localized part of the wave-field at the inclusion expressed as the total contribution from the cut-off frequency $\Omega_{\ast}$ for both integrals $I^{\mathrm{stop}}$ and $I^{\mathrm{pass}}$.
$A(w)=\frac{1}{\sqrt{\pi{\sqrt{1-w^{2}}}}}=\frac{1}{\sqrt{\pi}}+O(w^{2}),\quad w\to 0;\\ \qquad A(0)\neq 0.$ (3.26)
Though the last formula can be formally derived from Eq. (3.21), this way is incorrect since Eq. (3.21) is obtained assuming Eq. (3.2). The correct way to obtain formula (3.26) is to use the exact solution (2.9) Gavrilov2022ijhmt . Thus, the anti-localization in the linear system is an asymptotic weakening of the propagating component of the wave-field. The order of decay for the vanishing component is $1/\sqrt{t}$ in the uniform system and $1/\sqrt{t^{3}}$ in the system with a defect (near the defect).
@@
Now let us consider how a nonlinearity influences the anti-localization. To do this, following to Shishkina2023cmat , we proceed with numerical integration of $12n_{0}+$ 1 ODE (2.1), (2.4) with $n\in-n_{0}\dots n_{0}$, $n_{0}>\lambda t$, $\lambda=\operatorname{const}>1$. At first, consider the case of a heavy defect, where the impurity-induced localized mode does not exist for the corresponding linear system with $\beta_{n}=0$, and, hence, $v_{n}^{\mathrm{stop}}=0$.
@@
In Fig. 2, the numerically found particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ is shown for a fixed value of time. For reference, the wave-field in the corresponding linear system with $\beta_{n}=0$ given by Eqs. (3.6), (3.13)–(3.23), the wave-field in the uniform linear chain given by Eq. (2.9), and the numerically found wave-field in the corresponding uniform $\beta$-FPUT chain are also displayed. One can compare Fig. 2 with Fig. 2a in Shishkina2023cmat , where it is demonstrated that the numeric results for the linear system are in excellent agreement with asymptotic description Eqs. (3.6), (3.13)–(3.23), everywhere excepting the leading wave-fronts $n\simeq t$. One can see that nonlinearity essentially affects the wave-fields near the leading fronts $n\simeq t$, see Li2010 . However, in a central zone far from the leading fronts, the presence of nonlinearity does not alter the qualitative properties of the wave-fields essentially. Zeroing of the oscillation amplitude near $n=0$, i.e., the anti-localization, is observed in both linear and nonlinear systems with the defect. On the other hand, the anti-localization is not observed in both uniform systems under investigation.
@@
In Fig. 3a the numerically found solution for the particle velocity $v_{0}(t)$ versus the time $t$ is shown for the nonlinear system with the heavy defect. For reference, the same particle velocity in the corresponding linear system with $\beta_{n}=0$ given by Eq. (3.25) is also displayed. According to Fig. 3a, the order of decay for oscillation in the nonlinear system can be estimated the same as it is observed in the corresponding linear system, i.e., $t^{-3/2}$, see Eq. (3.25). The essential peculiarity of the oscillation in the nonlinear system with the defect, which can be observed in Fig. 3, is the oscillation asymmetry. Indeed, one can see that there exists a vanishing (apparently with the same power law $t^{-3/2}$) non-negative component of the particle velocity $v_{0}(t)$. Such a component does not exist in the corresponding linear system.
@@
In Fig. 3b the numerically found solution for the particle velocity $v_{0}(t)$ versus the time $t$ is shown for the nonlinear system without the defect. For reference, the same particle velocity in the corresponding linear system with $\beta_{n}=0$ given by Eq. (2.9), is also displayed. The order of decay for oscillation in the nonlinear system can be estimated as $t^{-1/2}$ as it is observed according to Eqs. (3.21), (3.22), (3.26) in the uniform linear system. Thus, there is no anti-localization in both nonlinear and linear systems without a defect.

### v2

@@ Abstract
Recently, a new general wave phenomenon, namely ‘‘the anti-localization of non-stationary linear waves’’, has been introduced and discussed (Shishkina et al., J. Sound. Vib. 553, 2023, 117673). This is zeroing of the propagating component for a non-stationary wave-field near a defect in infinitely long wave-guides. The phenomenon is known to be observed in both continuum and discrete mechanical systems with a defect, provided that the frequency spectrum for the corresponding homogeneous system possesses a stop-band. In this paper, we show that the anti-localization is also quite common for nonlinear systems. To demonstrate this, we numerically solve several non-stationary problems for an infinite strongly nonlinear $\beta$-FPUT chain with a defect. In our opinion, the anti-localization essentially influences the processes of heat transfer in linear and nonlinear lattices.
I Introduction
@@
In infinitely long uniform linear continuum and discrete systems, the perturbations caused by a pulse point load travel as waves  [^1] For a discrete system, it is more correct to speak about quasi-waves. along a system from the load application point to infinity. The presence of an inhomogeneity (or a defect) at the load application point can lead to accumulation or decumulation of the wave energy near the defect compared to the uniform case. Apparently, the first study where this fact was demonstrated is the classical paper by Lamb Lamb (1900), where it was shown that attaching a discrete mass-spring system into a continuum medium described by the 1D wave equation leads to the leakage of energy of a mass-spring system. On the other hand, such an attaching leads to the accumulation of energy in the wave-guide near the inhomogeneity, since in a uniform medium, described by the 1D wave equation, the perturbations caused by a point pulse load immediately leave the neighborhood of the application point, leaving it in the rest  [^2] This is due to the structure of the fundamental solution for the 1D wave equation Vladimirov (1971); Watanabe (2015)..
@@
In uniform linear dispersive systems, the amplitude of perturbations caused by a point pulse load vanishes with time not immediately but in a continuous manner. Generally, in 1D systems the order of decay for perturbations (e.g., for the particle velocities, the strains, or the displacements if the system is fixed somewhere and does not have rigid modes of motion) is $1/\sqrt{t}$, where $t\to\infty$ is time. This result is well-known; see, e.g., Whitham (1999); Slepyan (1972), and can be obtained by the asymptotic method of stationary phase applied to a representation of the exact non-stationary solution in the form of a Fourier integral. Again, introducing the inhomogeneity can lead to the accumulation or decumulation of the wave energy nearby.
@@
The effect of the accumulation can be related to the so-called trapped modes Ursell (1951); Indeitsev et al. (2007); Kuznetsov et al. (2002); Andrianov et al. (2013); Gavrilov et al. (2024), which can exist in dispersive systems. In infinite discrete systems (lattices or chains), the analogous phenomenon is known as the impurity-induced localized modes Montroll and Potts (1955); Kivshar (1991); Kivshar et al. (1997); Mamalui et al. (1999). The existence of such modes leads to trapping of some portions of the wave energy of a non-stationary process forever near a load application point at an inhomogeneity Kaplunov (1986); Teramoto and Takeno (1960) (we call this the localization of non-stationary waves). The whole frequency spectrum of our system can have a stop-band, where complex wave-numbers correspond to real frequencies. In such systems, so-called strong type of localization Luongo (2001, 1992) is quite common, which is related to a pole in the denominator of the system Green function in the frequency domain lying inside the stop-band. The rest of the energy (or all the wave energy if there is no localization) still travels to infinity as a vanishing propagating component of the wave-field, where perturbations again have the order of decay $1/\sqrt{t}$ (in a 1D system).
@@
Recently Shishkina and Gavrilov (2023a); Shishkina et al. (2023); Gavrilov et al. (2024), an opposite general wave phenomenon, which has been called the anti-localization of non-stationary waves, related to the decumulation of the energy near an inhomogeneity, has been introduced. It is observed in both discrete and continuum infinite linear systems, having the dispersion graph with a stop-band  [^3] The frequency spectrum of a discrete linear infinite system always has a stop-band Kossevich (1999) for high values of the frequency.. The presence of a stop-band leads to the existing of the cut-off frequency to which zero group velocity corresponds. The oscillation corresponding to the cut-off frequency accumulates near a point pulse source Hemmer (1959); Slepyan and Tsareva (1987); this can cause the phenomenon of the anti-localization. The existence of such a phenomenon is in agreement with some particular results previously obtained Kaplunov (1986); Hemmer (1959); Müller (1962); Müller and Weiss (2012); Kashiwamura (1962); Rubin (1963); Jackson (1978) for some systems, where the strong localization can take place. The anti-localization is zeroing of the leading-order term of order $1/\sqrt{t}$ in asymptotic expansions for perturbations at an inhomogeneity. Consequently, the next term of order $1/\sqrt{t^{3}}$ becomes a leading-order term there. Accordingly, the propagating perturbations, in an expanding with time neighborhood of the inhomogeneity, become asymptotically small compared to ones in zones closer to the leading wave-fronts. The amplitude of oscillation in the propagating component has a clear almost zero minimum at the inhomogeneity. The anti-localization can be observed in the absence of the localization as well as co-exist with the localization. In the latter case, after some time, a peak of the amplitude for oscillation also becomes clearly recognizable, being observed in the center of a zone where the perturbation amplitude is almost zero. Note that the term ‘‘weak anti-localization’’ is commonly known in modern physics Bergmann (1982). The anti-localization of non-stationary waves and the weak anti-localization are different phenomena. The weak anti-localization is related to the weak localization, whereas the anti-localization of non-stationary waves is observed in systems where the strong localization can be possible Shishkina et al. (2023).
@@
In the paper, we show that the anti-localization is not a specific peculiarity of linear systems. We demonstrate that it is also common for nonlinear ones. To achieve this, we solve several non-stationary problems for an infinitely long $\beta$-FPUT chain, introduced in the famous study Fermi et al. (1955), with a defect. In the strongly nonlinear case, we have no possibility to investigate the anti-localization in an analytic way using the method of stationary phase. Accordingly, we use a pure numerical approach, considering the previously obtained solutions for the corresponding linear systems as the reference ones. All numerical calculations presented in the paper have been performed twice using completely independent code and different numerical approaches. The first approach is to use routine scipy.integrate.odeint included in python framework scipy Virtanen et al. (2020), which is actually based on lsoda from the fortran library odepack. The second approach is related to the self-written code utilizing the semi-implicit Euler method Skeel and Cieśliński (2020). The results obtained by both approaches seem almost identical. The choice of a discrete system (not a continuum one) as the system under investigation is caused by the considerations of the simplicity of numerical solving.
@@
In our opinion, the anti-localization essentially influences the processes of heat transfer in linear and nonlinear lattices. In particular, the anti-localization can explain the occurrence of the Kapitza thermal resistance Gavrilov and Shishkina (2024); Gendelman and Paul (2021). The emergence of cold points at the defects demonstrated in linear and nonlinear cases Gendelman and Paul (2021), apparently, can be associated with this phenomenon, though the mathematical formulation used in Gendelman and Paul (2021) is different. The anti-localization can prevent the direct measurement of nonequilibrium temperature of a lattice as discussed in Hatano and Jou (2003), since the attaching the thermometer to the lattice leads to the distortion of the temperature field due to the anti-localization. The existence of the anti-localization makes a single-mode ansatz very effective Gavrilov et al. (2024) when looking for approximate non-stationary solutions describing non-stationary oscillation in the systems where the strong localization occurs.
@@
The structure of the paper is as follows. In Sect. II, we discuss the mathematical formulation of the problem in the dimensionless form. In Sects. IIIVI, several nonlinear problems, where we expect the emergence of the anti-localization, are investigated. In Conclusion (Sect. VII), we discuss the basic results of the paper. In Appendix A, we provide the procedure of the non-dimensionalization for the problem. In Appendix B, we reproduce some important formulas from Shishkina and Gavrilov (2023a) to explain to the reader what the anti-localization of non-stationary waves is in the linear case. In Appendix C, we provide some reference formulas for the semi-infinite linear chain.
II The mathematical formulation
@@
Consider an infinite ordered chain of point particles with nearest-neighbor interactions. All particles have an equal mass, whereas one particle has an alternated mass. The particles are labeled by the discrete variable $n\in\mathbb{Z}$, where $\mathbb{Z}$ is the set of all integers. Without loss of generality, we assume that the alternated mass corresponds to $n=0$. The particles are connected by nonlinear springs (i.e., by the bonds) with cubic nonlinearity also labeled by $n\in\mathbb{Z}$. We assume that a spring label equals the label of the preceding particle. All springs have equal stiffness excepting two of them with numbers n=1n=-1 and $n=0$, which connect the particle with $n=0$ and its neighbors. Thus, we consider a $\beta$-FPUT chain with a point impurity (or a defect) in the form of a mass-spring inclusion. Everywhere through the paper, excepting Sect. VI, we assume that the inclusion is symmetric  [^4] In Sect. VI we, in particular, consider a linear chain with an inclusion in the form of a nonlinear spring with the linear stiffness being equal to the spring stiffness in the rest of the chain.. The schematic of the system is presented in Fig. 1.
@@
Here, overdot denotes the derivative with respect to the dimensionless time $t$; $u_{n}(t)$ and $v_{n}(t)$ are the dimensionless displacement and particle velocity, respectively, of the particle with a label $n\in\mathbb{Z}$; $\beta$ is the dimensionless cubic stiffness for the bonds outside the defect; $m>0$ is the dimensionless mass of the particle at the defect; $K$ and $B$ are the linear and cubic dimensionless stiffnesses, respectively, for the bonds at the defect; $\delta_{n}$ is the Kronecker delta ($1$ if and only if $n=0$, 0 otherwise, $n\in\mathbb{Z}$).
The initial conditions are:
$u_{n}(0)=0,\qquad v_{n}(0)=m_{N}^{-1}\delta_{n-N}.$ (6)
Initial conditions in the form of Eq. (6) correspond to the pulse loading $\delta(t)$ applied to the particle with label $N$. Here $\delta(t)$ is the Dirac delta-function.
Note that we consider the strongly nonlinear system, i.e., the dimensionless parameter
$\bar{\beta}=\max_{n}\big{(}|\beta_{n}|\big{)}$ (7)
is assumed to be of order one, i.e., not a small quantity. Thus, we expect that the clear difference between a nonlinear solution and the corresponding linear one can be observed for finite values of time $t$.
@@
$m=1,\qquad K=1,\qquad B_{L}=B_{R}=\beta=0$ (8)
of an infinite uniform linear chain, the exact solution of the problem defined by Eqs. (1)–(6) for the particle velocities $v_{n}(t)$ is $V_{n-N}$ Schrödinger (1914); Hamilton (1940); Havelock (1910); Mühlich et al. (2020):
$V_{n}=J_{2n}(2t)=J_{2|n|}(2t),$ (9)
@@
III An infinite $\beta$-FPUT chain with an isotopic defect
@@
In the linear case $\beta=0$, this problem is considered in studies Rubin (1963); Shishkina and Gavrilov (2023a). In Shishkina and Gavrilov (2023a) an approximate solution of asymptotic nature, describing the wave-field as a whole and valid for large values of time $t$, is obtained. We reproduce in Appendix B some important formulas from Shishkina and Gavrilov (2023a) to explain to the reader what the anti-localization of non-stationary linear waves is.
@@
Now let us consider how a nonlinearity influences the anti-localization. To do this, following to Shishkina and Gavrilov (2023a), we proceed with numerical integration of $2n_{0}+1$ ODE (1), (4) with $n\in-n_{0}\dots n_{0}$, $n_{0}>\lambda t$, $\lambda=\operatorname{const}>1$. At first, consider the case of a heavy defect, where the impurity-induced localized mode does not exist for the corresponding linear system with $\beta_{n}=0$, and, hence, $v_{n}^{\mathrm{stop}}=0$.
@@
In Fig. 2, the numerically found particle velocity $v_{n}(t)$ versus the spatial variable $n\in\mathbb{Z}$ is shown for a fixed value of time. For reference, the wave-field in the corresponding linear system with $\beta_{n}=0$ given by Eqs. (24), (31)–(41), the wave-field in the uniform linear chain given by Eq. (9), and the numerically found wave-field in the corresponding uniform $\beta$-FPUT chain are also displayed. One can compare Fig. 2 with Fig. 2a in Shishkina and Gavrilov (2023a), where it is demonstrated that the numeric results for the linear system are in excellent agreement with asymptotic description Eqs. (24), (31)–(41), everywhere excepting the leading wave-fronts $n\simeq t$. The nonlinearity essentially affects the wave-fields near the leading fronts $n\simeq t$, see Li et al. (2010). In a central zone far from the leading fronts, we observe that the nonlinearity quantitatively alters the solution: the amplitude of oscillation in the nonlinear case is approximately a half of the amplitude in the linear case. Zeroing of the oscillation amplitude near $n=0$, i.e., the anti-localization, is observed in both linear and nonlinear systems with the defect. On the other hand, the anti-localization is not observed in both uniform systems under investigation.
@@
In Fig. 3a the numerically found solution for the particle velocity $v_{0}(t)$ versus the time $t$ is shown for the nonlinear system with the heavy defect. For reference, the same particle velocity in the corresponding linear system with $\beta_{n}=0$, i.e., the asymptotics given by Eq. (43) and the numerical solution, are also displayed. One can see that in the linear system the numerical solution very often converges to the asymptotics. On the other hand, we clearly deal with an essentially nonlinear regime, since the solutions in linear and nonlinear cases are not close to each other for finite times. According to Fig. 3a, the order of decay for oscillation in the nonlinear system can be estimated the same as it is observed in the corresponding linear system, i.e., $t^{-3/2}$, see Eq. (43). The essential peculiarity of the oscillation in the nonlinear system with the defect, which can be observed in Fig. 3, is the oscillation asymmetry. Indeed, one can see that there exists a vanishing (apparently with the same power law $t^{-3/2}$) non-negative component of the particle velocity $v_{0}(t)$. Such a component does not exist in the corresponding linear system Shishkina and Gavrilov (2023a).
@@
In Fig. 3b the numerically found solution for the particle velocity $v_{0}(t)$ versus the time $t$ is shown for the nonlinear system without the defect. For reference, the same particle velocity in the corresponding linear system with $\beta_{n}=0$ given by Eq. (9), is also displayed. The order of decay for oscillation in the nonlinear system can be estimated as $t^{-1/2}$ as it is observed according to Eqs. (39), (40), (44) in the uniform linear system. Thus, there is no anti-localization in both nonlinear and linear systems without a defect.